# Unifying Specialized Visual Encoders for Video Language Models

### v1

@@ 4.1 Comparison to State of the Art
Table 1: Comparison of different multimodal LLMs on video reasoning benchmarks. We employ ChatGPT to evaluate performance following Video-ChatGPT where applicable222 Video-ChatGPT’s and Video-LLaVA’s author-reported numbers on TGIF are incomparable as they were on a subset of the dataset. See https://github.com/PKU-YuanGroup/Video-LLaVA/issues/37. (version gpt-3.5-turbo-0613). * denotes our evaluation of using the author provided checkpoint. The first five datasets were used as development sets; the last three were held-out for our final evaluation.
Table 1 tabulates the performance of MERV (frozen) and (full). We compare our model to the existing state-of-the-art works, including Video-LLaVA (Lin et al., 2023a) that share our training data mixture, and other VideoLLMs (Li et al., 2023c; e; Maaz et al., 2023; Yu et al., 2024; Zhang et al., 2023a; b). We find that our method, generating video representations using multiple visual encoders that specialize in different skills of video understanding, outperforms Video-LLaVA across nearly all of the benchmarks, with a 3.2% gain on MSVD and a 3.7% gain on ActivityNet. Both of our methods perform better overall than Video-LLaVA, even when using less data with just Stage 2 as shown by the MERV numbers. While MERV (full) is not a strict improvement to MERV, it still improves on some difficult benchmarks with its additional video-language alignment. We believe that outside of these testing sets, MERV (full) is a better model overall and recommend using this recipe when possible. Compared to LLaMA-VID (Li et al., 2023e), which uses a different training mix, we also better in nearly all benchmarks, up to around 4.5% across Perception Test, ActivityNet, and TVQA. MERV (full) outperforms the previous state-of-the-art on the Perception Test zero-shot with 48.4%, compared to SeViLa (Yu et al., 2024) with a 46.2% accuracy. Overall, our design shows a significant improvement over Video-LLaVA and prior methods as a whole.
4.2 Ablations
In this section, we justify the design choices for our architecture, covering our projectors, feature fusion strategies, and training recipes. Our ablations are done with the MERV (frozen) recipe.
4.2.1 Pre-Fusion Projectors
The first module we investigate is our projectors, which serve to connect each encoder from its pretrained embedding space to a common embedding space. We test two types of projectors: image-level, which operate on frames independently, and video-level, which aggregate information across frames. Projector details are provided in Section B.1 in the Appendix. We report average performance across our development sets of MSVD, MSRVTT, TGIF, and Perception Test.
Pre-fusion projector. Table 2a tabulates our projector ablation on average accuracy, parameter count, and FLOPs, with the default settings of LanguageBind as the single vision encoder and an 8 frame 64 token projection output. We find that 2D average pooling is the best, with 55.86 average accuracy, even better than using no projector on the full 257 token representation (as in Video-LLaVA (Lin et al., 2023a)), with 54.76 average accuracy. It also has no trainable parameters and the fewest FLOPs. The projection serves as a form of feature selection, allowing the LLM to efficiently reason only about the most relevant information. One result worth noting is the poor performance (52.12, 3.7% lower) of the attentive resamplers, a popular projector choice. They are agnostic to structure, which leads them to being weaker projectors for us. However, it would likely be better with higher quality data from which the transformer can learn. Increasing the frame resolution from 8 to 16 was also a large improvement, showing that increasing temporal resolution is still important. This highlights the importance of aligning representations with their spatial and temporal structure, especially for video models, which extract many more frames of visual information.
Projector token length. Similarly, we ablate the output token length of the projector. Table 2b tabulates the performance of producing 1 to 256 tokens for a 2D average pooling projector with 16 frames as input. We see that the performance peaks at 64 tokens with 69.08 MSVD accuracy, with worse performance for longer token lengths (max 68.65). This balances the number of tokens used for condensing the visual embedding while also minimizing the extra processing needed by the LLM.
4.2.2 Feature Fusion Strategies
@@
(c) Feature fusion strategy. Cross-Attn additive mixing is best overall among all the strategies on accuracy, for its FLOPs.
Table 2: Ablating design choices. We highlight our defaults in orange and bold the best results. Average accuracy is on MSVD, MSRVTT, TGIF, and Perception Test. Full results are in the Appendix.
@@
Next, we test different strategies for fusing the information from all of the features, with detailed breakdowns in Table 2c. First, we evaluate two popular concatenation methods, in either the token sequence dimension, or the channel dimension followed by an MLP projector for matching the LLM dimension. While sequence-wise concatenation is widely used in multimodal LLMs (Tong et al., 2024), our method outperforms it while using significantly less computation, with a 56.8% average accuracy compared to 54.4%, while also using 2.5$\times$ fewer FLOPs. Concatenation channel-wise reaches a similar performance of 56.6% and a lightweight cost. However, our cross-attention shows slightly better performance, with the additional benefit of having accessible encoder weightings for analysis, so we stick with cross-attention for our final design. We also try different methods of additive mixing as an ablation. The last two rows of Table 2c show the performance when either learning the additive weights directly as a learnable scalar or by fixing the weights to be 0.25 for each of 4 encoders. We see that using cross attention outperforms both methods by 1.8% and 2.6%, as our feature fusion module can dynamically generate better fused embeddings given the visual input.
@@ 4.2.3 Training Recipes
Finally, we also compare different training recipes based on the literature and our own expertise. Traditional rule-of-thumb follows that of the original LLaVA recipe: a Stage 1 pre-training on captioning data to align the projectors only, and a Stage 2 instruction tuning on multi-turn complex reasoning data for both projectors and the LLM. Many recent works have attempted some combination of other strategies, such as unfreezing the vision encoders in Diao et al. (2024) or skipping the Stage 1 in Karamcheti et al. (2024).
We systematically map out this landscape, fixing our dataset to be the same as Video-LLaVA’s. Unlike them, we found that the Stage 1 phase did not help much when training only the projectors and feature fusion (Table 5 in the Appendix), with roughly the same average accuracy. Stage 2 instruction tuning alone leads to similar results in 43% of the total time, so we adopt this recipe for efficiency and refer to it as MERV (frozen). This recipe is still unsatisfying as it leaves a large amount of data, approximately 1.3M vision-text pairs, unused for training. In our empirical observations, we found that video-language alignment was not very strong. The distributions of language used in video datasets and benchmarks sparsely overlap based on their sentence embeddings, which could be impacting our ability to generalize zero-shot on downstream benchmarks. We address this by unfreezing the LLM during Stage 1 to better learn this alignment, improving performance on a few key benchmarks, especially Perception Test, by up to 2.2%. We call this recipe MERV (full).
Figure 3: Extra encoders incur minimal step time overhead. Here we add encoders in the order of DINOv2, LanguageBind, SigLIP, ViViT, plotted alongside the slowest single encoder in each group.
As another ablation, we train MERV on a single stage comprised of the Stage 1 and Stage 2 data mixed together (bottom of Appendix Table 5). Surprisingly, this does worse than the explicit two stage training recipe. We attribute this to the explicit types of data in each stage being a form curriculum learning, showing that these stages are still important for optimal performance.
Finally, we provide evidence for the efficiency of our method (see Figure 3). We use the default FSDP sharding strategy PyTorch provides; it is not currently possible to specify explicit plans for which modules go where (but may be possible as FSDP matures). However, even with this basic strategy, our method is dominated by the slowest single encoder present, incurring very little additional overhead from extra encoders due to this parallelization, making it cost-efficient to scale up in the number of encoders.
@@ 4.3 Ensemble Composition
Can we make use of more encoders? The conventional wisdom is to use a single encoder, typically a contrastively trained vision-language model like CLIP, SigLIP, or LanguageBind (Radford et al., 2021; Zhai et al., 2023; Zhu et al., 2023), in a VideoLLM. In Figure 4a, we show the four single encoder models corresponding to each of our chosen encoders using their full embeddings. They not only all perform worse than MERV but also use more FLOPs, as without our pre-fusion projectors, their sequence lengths are at least 4×4$\times$4 × ours.
Are each of the encoders contributing? To affirm that this set of four encoders is actually beneficial for improving understanding, we train three-encoder VideoLLMs under the same strategy, but removing a different encoder each time. Each of these models does worse based on the strength of the encoder removed, meaning that MERV is using their knowledge (Fig. 4a). The minor drop in FLOPs illustrates how most of the computation is still dominated by the LLM, not the vision encoders.
@@ 5 Analysis
(a) Visual Encoder Subsets. MERV outperforms single-encoder VideoLLMs, with our feature projectors unlocking more computational effiency. Removing any encoder also reduces MERV performance.
(b) SSv2-MCQ and Temporal. Temporal denotes performance on 12 selected classes where actions are indistinguishable if played in reverse.
Figure 4: Analysis plots supporting our design of multiple encoders, from their accuracy to their skill specializations. Average accuracy is across MSVD, MSRVTT, TGIF, and Perception Test. Full results are in the Appendix Tables 7810.
@@ 5.1 Cross Attention Weights Activate On Corresponding Videos
We first understand our method by looking at the cross attention weights on our 4 benchmark datasets (MSRVTT, TGIF, MSVD, and Perception Test), and visualizing the videos which have the highest attention weight for each of the encoders on Figure 12 in the Appendix. This lets us see what types of videos activate each encoder the most. As expected, ViViT attention weights are highest on videos with large motion, as ViViT has strong temporal understanding. Meanwhile, SigLIP is utilized for videos that have textual data in the video, likely due to being vision-language contrastively trained, especially with textual data during training. DINOv2 and LanguageBind are both preferred by videos with static scenes, but LanguageBind is preferred for videos with foreground motion.
@@ 5.2 MERV can capture visual skills of different encoders
Next, we ask if our model effectively captures knowledge from its encoders. We first answer through our previous open-ended QA benchmarks. To assess the performance across different visual tasks, we create “pseudo”-skill categories by looking at the first word of the question sentence, which are often WH-words. They can be viewed as a proxy of skills required to solve the task. For example, Where requires spatial understanding and When requires temporal understanding. Figure 5 shows the relative performance of different visual encoders. While the contrastive models generally dominate each category, no single encoder performs best in all tasks. LanguageBind, for example, performs the best in TGIF-What with 46.23%, while DINOv2 performs on par with the best in MSVD-Who with 82.12%. Our method combines different encoders into an unified representation and consistently matches or improves the best-performing encoder. Raw numbers are in Table 9 in the Appendix.
@@ 5.3 MERV Can Intuit Motion and General Understanding Simultaneously
We take another angle to quantifying how well our model learns from each of its individual encoders by looking back to classic video action recognition datasets. They are commonly used for general action understanding, but they can also be used to evaluate finer grained capabilities. We are most interested in both general understanding and distinguishing actions which are temporally ambiguous, i.e. indistinguishable when reversed in time, such as Pulling [something] from left to right and Pulling [something] from right to left. This offers a fair analysis of both general and video understanding.
We turn to the Something-Something v2 (Goyal et al., 2017) (SSv2) dataset, where the goal is to classify an input video into one of 174 classes, e.g., Pulling [something] from left to right. This allows us to analyze our model’s understanding of temporal-spatial interaction with minimal distractions from scene understanding and real-world semantics. However, evaluating SSv2 as a zero-shot open-ended task is difficult with a long tail of specific categories. Thus we reformat the dataset into 5-way multiple-choice questions (MCQ) and fix the prompt to be “How is the object in the video being interacted with?”. Incorrect choices were randomly sampled from the other 173 classes. We call this benchmark “SSv2 - MCQ” to distinguish it from the original classification task. To form the temporal subset, we selected 12 classes a priori from SSv2, where the action is indistinguishable if reversed in time, e.g., Pulling [something] from left to right and Pulling [something] from right to left.
Figure 4b plots the performance of MERV and single-encoder models on this temporal subset ($x$-axis) against the full dataset ($y$-axis). We see that ViViT, which often falls short in other Video QA benchmarks, surprisingly performs better than other encoders at 39.77%, which is 9.19% higher than the next closest model LanguageBind. However for the full dataset, ViViT suffers with a worse performance of 26.78%, as ViViT’s strength is on temporal understanding despite lacking in vision-language understanding. Contrastive encoders have the upper-hand in most other classes.
@@
Figure 6: Single-Encoder Performance Difference in Something-Something v2 - MCQ. ViViT shows better performance on tasks where temporal understanding is crucial, while LanguageBind and SigLIP show better performance where task can be solved from single-frame understanding.
We believe that the architecture, datasets, and objective of each model causes these difference. ViViT processes spatial-temporal tubelets for embeddings, leading to better temporal understanding despite only being pre-trained on Kinetics-400 classification. SigLIP uses image-based ViT with no temporal layer has limited temporal understanding, but has a greater knowledge due to its larger training set and contrastive objective. MERV, at 42%, shows better performance compared to all these single-encoder models via leveraging strength of all the individual encoders. MERV (full) performs better than both VideoLLaVA (Lin et al., 2023a) and the 7B and 13B variants of LLaMA-Vid (Li et al., 2023e).
Finally, we also plot the performances of 10 SSv2 classes where the performance difference between ViViT and SigLIP is largest in Figure 6. We see that actions which cannot be inferred from a single frame are the ones that ViViT performs better, e.g., Moving [something] down is indistinguishable from Moving [something] up if temporal information is omitted. Meanwhile, SigLIP performs better for classes where understanding the semantics of the scene can hint the action that is happening, e.g., if the video contains a cup and a bottle of water, one can easily expect Showing [something] is empty without watching the full video. See Appendix Figure 9 for sample videos of the 10 classes.

### v2

@@ 3.2 Spatio-Temporally Aligned Representations for Feature Fusion
3.3 Implementation Efficiencies
Parallelized visual encoding. At a first glance, using multiple encoders seems to be a large cost to pay when comparing the raw FLOPs and parameters. However, a key benefit of the LLaVA style architecture is that the entire feature extraction and projection pipeline can happen in parallel. To make this possible, we build on top of recent powerful advances in parallel processing for LLMs and use PyTorch’s Fully Sharded Data Parallel (Zhao et al., 2023). As the video encoders themselves are much smaller than the LLM blocks, and their visual encoding processes are completed in roughly the same amount of time, most of the overhead in running four encoders is already covered by having just one encoder (ref. Figure 3). We provide some timing numbers in Section 4.2.3 and find that our step time is similar to that of the single-encoder methods.
Our code is built on top of the Prismatic VLM codebase (Karamcheti et al., 2024), which efficiently implements vision-language model (VLM) training. We add the ability to handle videos and an arbitrary number of visual encoders, along with many useful features for training. Our training is efficient for using multiple visual models, completing in under 24 hours using 8 L40-48GB GPUs, and down to 8 hours using 8 H100s. In contrast, the Video-LLaVA codebase runs the Stage-2 training in around 38 hours on the same L40 setup and could not easily support multiple encoders in our initial attempts.
MERV frozen and full. Many different recommendations for training LLaVA style models have been made since its inception. This is only made more complicated by the introduction of new datasets with every new VideoLLM architecture, making it difficult to properly determine the best recipe for one’s own setup. We intentionally fix our dataset to be the same as Video-LLaVA’s so we can isolate the impacts of the training setup, from which we find two viable settings: MERV (frozen) and MERV (full). The original Video-LLaVA’s recipe has two stages: a Stage 1 pre-training on captioning data to align only the connector between the pre-trained vision encoder and LLM, and a Stage 2 instruction tuning on both connector and the LLM.
  • MERV (frozen), which performs only the Stage 2 instruction tuning inspired by Karamcheti et al. (2024); it achieves similar results to the original Video-LLaVA recipe in only 43% of the time;
  • MERV (full), which undergoes both of the Stage 1 captioning pre-training and the Stage 2 instruction tuning, but also unfreezes the LLM during Stage 1 for a slight improvement on a few benchmarks.
As MERV (frozen) is faster to train with similar performance, we adopt that recipe by default for analysis, and interchangeably use MERV to refer to it for simplicity from here on out. Detailed analysis is provided in Section 4.2.3.
4 Experimental Results
In this section, we show that our method outperforms prior works across standard video-language benchmarks before moving onto an in-depth analysis of our method and its specializations in the next section.
Datasets and training procedure details. For fair comparison, our data mix is the same as Video-LLaVA (Lin et al., 2024). The Stage 1 data is single-turn concise captioning, with 558k (image, text) pairs from LAION filtered by LLaVA (Liu et al., 2023) and 702k (video, text) pairs from Valley (Luo et al., 2023). The Stage 2 data is multi-turn conversations, detailed captioning and reasoning, with 665k (image, text) pairs from LLaVA and 100k (video, text) instructions from Video-ChatGPT (Maaz et al., 2024).
Table 2: Ablating design choices. We highlight our defaults in orange and bold the best results. Average accuracy is on MSVD, MSRVTT, TGIF, and Perception Test. Full metric results are in the Appendix.
@@ 4 Experimental Results
(c) Feature fusion strategy. Cross-Attn additive mixing is the best overall among all the strategies on accuracy, for its FLOPs.
4.1 Comparison to State of the Art
Table 1 tabulates the performance of MERV (frozen) and (full). We compare our model to the existing works, including Video-LLaVA (Lin et al., 2024) that share our training data mixture, and other VideoLLMs (Li et al., 2023c, 2024b; Maaz et al., 2024; Yu et al., 2024b; Zhang et al., 2023, 2024b). We find that our method, generating video representations using multiple visual encoders that specialize in different skills of video understanding, outperforms Video-LLaVA across nearly all of the benchmarks, with a 4.1% gain on Perception Test, a 3.7% gain on ActivityNet and a 4.6% gain on TVQA. Both of our methods perform better overall than Video-LLaVA, even when using less data with just Stage 2 as shown by the MERV (frozen) numbers. While MERV (full) is not a strict improvement to MERV, it still improves on some difficult benchmarks (e.g., Perception Test) with its additional video-language alignment. We believe that this makes MERV (full) generalize better in unseen settings outside of these testing benchmarks, and recommend using this recipe when possible. Compared to LLaMA-VID-7B, which uses a different training mix, we are better in nearly all benchmarks, up to around 4.5% across Perception Test, ActivityNet, and TVQA. Moreover, MERV (full) outperforms SeViLA on the Perception Test zero-shot with 48.4%, compared to 46.2%. Overall, our design shows a significant improvement over Video-LLaVA and prior methods as a whole.
4.2 Ablations
In this section, we justify our design choices for the projectors, feature fusion strategies, and training recipes. Our ablations are done with the MERV (frozen) recipe.
4.2.1 Pre-Fusion Projectors
The first module we investigate is our projectors, which serve to connect each encoder from its pre-trained embedding space to a common embedding space. We test two types of projectors: image-level, which operate on frames independently, and video-level, which aggregate information across frames. Projector details are provided in Section A.3 in the Appendix. We report average performance across our development sets of MSVD, MSRVTT, TGIF, and Perception Test.
Pre-fusion projector. Table 2a tabulates our projector ablation on average accuracy, parameter count, and FLOPs, with the default settings of LanguageBind as the single vision encoder and an 8 frame 64 token projection output. We find that 2D average pooling is the best, with 55.86% average accuracy, even better than using no projector on the full 257 token representation (as in Video-LLaVA (Lin et al., 2024)), with 54.76% average accuracy. It also has no trainable parameters and the fewest FLOPs. The projection serves as a form of feature selection, allowing the LLM to efficiently reason only about the most relevant information.
One result worth noting is the poor performance (52.12%, 3.7% lower) of the attentive resampler, a popular projector choice. The attentive resampler is less responsive to the input spatio-temporal structure compared to other projectors, which leads it to being a weaker projector for us. However, it would likely be better with higher quality data from which the transformer can learn. Increasing the frame resolution from 8 to 16 was also a large improvement, showing that increasing temporal resolution is still important. This highlights the importance of aligning representations with their spatial and temporal structure, especially for video models, which extract many more frames of visual information.
Projector token length. Similarly, we ablate the output token length of the projector. Table 2b tabulates the performance of producing 1 to 256 tokens for a 2D average pooling projector with 16 frames as input. We see that the performance peaks at 64 tokens with 69.08% MSVD accuracy, with worse performance for longer token lengths (max. 68.65%). This balances the number of tokens used for condensing the visual embedding while also minimizing the extra processing needed by the LLM.
4.2.2 Feature Fusion Strategies
@@
Next, we test different strategies for fusing the information from all of the features, with detailed breakdowns in Table 2c. First, we evaluate two popular concatenation methods, in either the token sequence dimension, or the channel dimension followed by an MLP for matching the LLM dimension. While sequence-wise concatenation is widely used in multimodal LLMs (Tong et al., 2024b), our method outperforms it while using significantly less computation, with a 56.8% average accuracy compared to 54.4%, while also using 2.5$\times$ fewer FLOPs. Concatenation channel-wise reaches a similar performance of 56.6% and a lightweight cost. However, our cross-attention shows better average performance, better or on par on 3 out of 4 benchmarks (ref. Appendix Table 6c), with the additional benefit of having accessible encoder weightings for analysis, so we stick with cross-attention for our final design. We also try different methods of additive mixing as an ablation. The last two rows of Table 2c show the performance when either learning the additive weights as a single scalar or by fixing the weights to be 0.25 for each of the 4 encoders. We see that using cross attention outperforms both methods by 1.8% and 2.6%, as our feature fusion module can dynamically generate better fused embeddings based on the visual input.
@@ 4.2.3 Training Recipes
Finally, we also compare different training recipes based on the literature and our own expertise. Apart from training recipes mentioned in Section 3.3, many recent works have attempted some combination of other strategies, such as unfreezing the vision encoders in Diao et al. (2024), which usually requires more training data.
We systematically map out this landscape, fixing our dataset to be the same as Video-LLaVA’s. Unlike them, we found that the Stage 1 phase did not help much when training only the projectors and feature fusion (ref. Appendix Table 5), with roughly the same average accuracy. Stage 2 instruction tuning alone leads to similar results in 43% of the total time, so we adopt this recipe for efficiency and refer to it as MERV (frozen). This recipe is still unsatisfying as it leaves a large amount of data, approximately 1.3M vision-text pairs, unused for training. In our empirical observations, we found that the resulting video-language alignment was suboptimal. The distributions of language used in video training datasets and benchmarks sparsely overlap based on their sentence embeddings, which could be impacting our ability to generalize zero-shot on downstream benchmarks. We address this by unfreezing the LLM during Stage 1 to better learn this alignment, improving performance on a few key benchmarks, especially Perception Test, by up to 2.2%. We call this recipe MERV (full).
Figure 3: Extra encoders incur minimal step time overhead. We add encoders in the order of DINOv2, LanguageBind, SigLIP, ViViT, plotted alongside the slowest single encoder in each group.
As another ablation, we train MERV on a single stage comprised of the Stage 1 and Stage 2 data mixed together (ref. bottom of Appendix Table 5). Surprisingly, this does worse than the explicit two stage training recipe. We attribute this to the explicit types of data in each stage being a form curriculum learning, showing that these stages are still important for optimal performance.
Finally, we provide evidence for the efficiency of our method (ref. Figure 3). We use the default FSDP sharding strategy PyTorch provides; it is not currently possible to specify explicit plans for which modules go where (but may be possible as FSDP matures). However, even with this basic strategy, our method is dominated by the slowest single encoder present, incurring very little additional overhead from extra encoders due to this parallelization, making it cost-efficient to scale up in the number of encoders.
@@ 4.3 Ensemble Composition
Can we make use of more encoders? The conventional wisdom is to use a single encoder, typically a contrastively trained vision-language model like CLIP, SigLIP, or LanguageBind (Radford et al., 2021; Zhai et al., 2023; Zhu et al., 2024a), in a VideoLLM. In Figure 4a ( ) , we show the four single encoder models corresponding to each of our chosen encoders using their full embeddings. They not only all perform worse than MERV but also use more FLOPs, as without our pre-fusion projectors, their sequence lengths are at least 4×4$\times$4 × ours.
Are each of the encoders contributing? To affirm that this set of four encoders is actually beneficial for improving understanding, we train three-encoder VideoLLMs under the same strategy, but removing a different encoder each time. Each of these models does worse based on the strength of the encoder removed, meaning that MERV is using their knowledge (ref. Fig. 4a, ). The minor drop in FLOPs illustrates how most of the computation is still dominated by the LLM, not the vision encoders.
@@ 5 Analysis
Figure 4: Analysis plots supporting our design of multiple encoders, from their accuracy to their skill specializations.
(a) Visual Encoder Subsets. MERV outperforms single-encoder VideoLLMs ( ), with our feature projectors unlocking more computational efficiency. Removing any encoder also reduces MERV performance ( ). Average accuracy is across MSVD, MSRVTT, TGIF, and Perception Test. (b) SSv2-MCQ and Temporal. Temporal denotes performance on 12 selected classes where actions are indistinguishable if played in reverse (ref.  Sec. 5.3). Full results are in the Appendix Tables 7810.
@@ 5.1 Cross Attention Weights Activate on Corresponding Videos
We first understand our method by looking at the cross attention weights on our 4 benchmark datasets (MSRVTT, TGIF, MSVD, and Perception Test), and visualizing the videos which have the highest attention weight for each of the encoders in Figure 5. This lets us see what types of videos activate each encoder the most. As expected, ViViT attention weights are highest on videos with large motion, as ViViT has strong temporal understanding. Meanwhile, SigLIP is utilized for videos that have textual data in the video, likely due to being vision-language contrastively trained, especially with textual data during training. DINOv2 and LanguageBind are both preferred by videos with static scenes, but LanguageBind is preferred for videos with foreground motion.
LanguageBind
DINOv2
ViViT
SigLIP
Figure 5: Videos that give the highest attention weight for each of the encoders. The right-most column shows the average frame of the video. For more examples, see Figure 11 in the Appendix.
@@ 5.2 MERV Can Capture Visual Skills of Different Encoders
Next, we ask if our model effectively captures knowledge from its encoders. We first answer through our previous open-ended QA benchmarks. To assess the performance across different visual tasks, we create “pseudo”-skill categories by looking at the first word of the question sentence, which are often WH-words. They can be viewed as a proxy of skills required to solve the task. For example, Where requires spatial understanding and When requires temporal understanding. Figure 6 shows the relative performance of different visual encoders. While the contrastive models generally dominate each category, no single encoder performs best in all tasks. LanguageBind, for example, performs the best in TGIF-What with 46.23%, while DINOv2 performs on par with the best in MSVD-Who with 82.12%. Our method combines different encoders into an unified representation and consistently matches or improves the best-performing single-encoder model. Raw numbers are in Table 9 in the Appendix.
@@ 5.3 MERV Can Intuit Motion and General Understanding Simultaneously
We take another angle to quantifying how well our model learns from each of its individual encoders by looking back to classic video action recognition datasets. They are commonly used for general action understanding, but they can also be used to evaluate finer-grained capabilities. We are most interested in both general understanding and distinguishing actions which are temporally ambiguous, i.e., indistinguishable when reversed in time, such as Pulling [something] from left to right and Pulling [something] from right to left. This offers a fair analysis of both general and finer-grained video understanding.
We turn to the Something-Something v2 (Goyal et al., 2017) (SSv2) dataset, where the goal is to classify an input video into one of 174 classes, e.g., Pulling [something] from left to right. This allows us to analyze our model’s understanding of temporal-spatial interaction with minimal distractions from scene understanding and real-world semantics. However, evaluating SSv2 as a zero-shot open-ended task is difficult with a long tail of specific categories. Thus we reformat the dataset into 5-way multiple-choice questions (MCQ) and fix the prompt to be “How is the object in the video being interacted with?”. Incorrect choices were randomly sampled from the other 173 classes. We call this benchmark “SSv2 - MCQ” to distinguish it from the original classification task. Additionally, we selected 12 classes a priori from SSv2, where the action is indistinguishable if reversed in time, e.g., Pulling [something] from left to right and Pulling [something] from right to left to form a subset that can test the model’s ability in temporal understanding.
Figure 4b plots the performance of MERV and single-encoder models on this temporal subset ($x$-axis) against the full dataset ($y$-axis). We see that the ViViT single-encoder model, which often falls short in other video QA benchmarks, surprisingly performs better than other encoders at 39.77% on temporal subset, which is 9.19% higher than the next closest model LanguageBind. However for the full SSv2 - MCQ, ViViT suffers with a worse performance of 26.78%, as ViViT’s strength is on temporal understanding despite lacking in vision-language understanding. Contrastive encoders have the upper-hand in most other classes.
@@
We believe that the architecture, datasets, and objective of each model causes these difference. ViViT processes spatial-temporal tubelets for embeddings, leading to better temporal understanding despite only being pre-trained on Kinetics-400 classification. SigLIP uses image-based ViT with no temporal layer has limited temporal understanding, but has a greater knowledge due to its larger training set and contrastive objective. MERV, at 42%, shows better performance compared to all these single-encoder models via leveraging strength of all the individual encoders. MERV (full) performs better than both VideoLLaVA (Lin et al., 2024) and the 7B and 13B variants of LLaMA-Vid (Li et al., 2024b).
Figure 7: Single-Encoder Performance Difference in Something-Something v2 - MCQ. ViViT shows better performance on tasks where temporal understanding is crucial, while LanguageBind (LB) and SigLIP show better performance where task can be solved from single-frame understanding.
Finally, we also plot the performances of six SSv2 classes where the performance difference between ViViT and SigLIP is largest in Figure 7. We see that actions which cannot be inferred from a single frame are the ones that ViViT performs better, e.g., Pushing [something] from left to right is indistinguishable from Pushing [something] from right to left if temporal information is omitted. Meanwhile, SigLIP performs better for classes where understanding the semantics of the scene can hint the action that is happening, e.g., if a frame where a blanket is dropped on an object is given, one can easily expect Covering [something] with [something] without watching the full video. See Appendix Figure 9 for sample videos of the these classes.